OpenStudy (anonymous):

3x+y+7+90=180 what is x and y

OpenStudy (anonymous):

so you want y as a function of x? then x is a function of y?

OpenStudy (anonymous):

3x+y+7+90=180 what is x and y separate y from the rest to find y for ex if you put y on the left handside its like y=?

OpenStudy (anonymous):

Just solve for y and x by isolating them: y=-3x+180-90-7 y=-3x+83 3x=-y+180-90-7 3x=-y+83 x=(-1/3)y+(83/3)